# A Tie-breaking based Local Search Algorithm for Stable Matching Problems

### v1

@@ 5.3.1 MAX-SMTI Problems
First, we compare the solution quality of different algorithms in solving small-sized MAX-SMTI problems, including 220 configurations. The maximum number of iterations is set to 3000 for TBLS and TBLS-E, and 50000 for LTIU, AS, and MCS, to fully leverage the capabilities of the latter three local search algorithms.
@@
Figure 1(a) displays the number of singles produced by various algorithms in solving small-sized MAX-SMTI problems. Our TBLS algorithm achieves the highest number of wins and the lowest average number of singles across 220 configurations. TBLS-E closely follows, also demonstrating excellent performance with an average number of singles very similar to that of TBLS. Furthermore, TBLS-E stands out as the top performer measured by SECost, as illustrated in Figure 1(b). In almost all configurations, TBLS-E achieves the fairest matchings, significantly outperforming other algorithms.
Next, we assess the solution quality in addressing large-scale MAX-SMTI problems across 110 configurations, maintaining the same settings for the maximum number of iterations as used for small-sized MAX-SMTI problems. Due to its slow performance, LTIU is excluded from this evaluation. As the instances become harder, the performance gap between algorithms widens. Regarding the number of singles, TBLS remains the winner in most configurations, with approximately 0.2 fewer singles than the second-place TBLS-E and about 1.5 fewer singles than the third-place GSA2, as shown in Figure 2(a). In terms of fairness, TBLS-E consistently leads in most configurations, with an average SECost approximately 100 lower than the second-place MCS, as illustrated in Figure 2(b).
@@
Finally, we evaluate the execution time of different local search algorithms in solving small-sized and large-sized MAX-SMTI problems. For these comparisons, the maximum number of iterations is set to 3000 for all the algorithms. Figure 3(a) illustrates the execution time for small-sized MAX-SMTI problems. Although all these algorithms, except LTIU, complete their searches within an average of 0.1 seconds, our TBLS algorithm performs the fastest in most configurations. This gap becomes more pronounced as the problem size increases, which is shown in Figure 3(b). For large-sized MAX-SMTI problems, TBLS and TBLS-E are the top performers in nearly all configurations. On average, the first-place TBLS and the second-place TBLS-E run about three times and 1.3 times faster than the third-place MCS, respectively.
(a) Small-sized MAX-SMTI - execution time
(b) Large-sized MAX-SMTI - execution time
@@ 5.3.2 MAX-HRT Problems
We begin by evaluating the solution quality of different algorithms for solving MAX-HRT problems of small and large sizes. The maximum number of iterations is still set to 3000 for TBLS. To maximize the performance of other local search algorithms, the maximum number of iterations is set to 5000 for both HR and MCA.
Figure 4(a) and Figure 4(b) illustrate the number of unassigned positions produced by various algorithms in addressing small-sized and large-sized MAX-HRT problems, respectively. Overall, TBLS outperforms all other algorithms in both problem scales, consistently yielding matchings with the fewest unassigned positions. Notably, for large-sized MAX-HRT problems, TBLS produces approximately 1.7 fewer unassigned positions than the second-place algorithm, HPA, and about 10.8 fewer unassigned positions than HR, the second-best among all local search algorithms.
@@
(a) Small-sized MAX-HRT - execution time
(b) Large-sized MAX-HRT - execution time
@@
After evaluating the solution quality, we shift our focus to analyzing execution time performance. The maximum number of iterations is set to 3000 for all the algorithms. Figure 5(a) and Figure 5(b) show the execution time for small-sized and large-sized MAX-HRT problems, respectively. In both problem scales, TBLS emerges as the algorithm with the highest number of wins, indicating it runs the fastest in a large number of configurations. In particular, for large-sized MAX-HRT problems, TBLS operates approximately three times faster than MCA and about four times faster than HR.
@@ 6 Conclusion
This paper proposed a tie-breaking based local search algorithm, named TBLS, to solve the MAX-SMTI and MAX-HRT problems. We also introduced an equity-focused variant, TBLS-E, designed to find relatively fair matchings for the MAX-SMTI problem. Experimental results demonstrate that TBLS consistently outperforms other algorithms in achieving larger matching sizes for both the MAX-SMTI and MAX-HRT problems. Additionally, for the MAX-SMTI problem, TBLS-E yields matchings with lower sex equality costs while preserving matching sizes comparable to those produced by TBLS. Both TBLS and TBLS-E exhibit faster execution time than other local search algorithms when solving large-scale problems. In the future, we plan to incorporate neural networks and reinforcement learning techniques, such as G3DQN [24] and NeuRewriter [5], to improve the evaluation function and the refinement process.

### v2

@@ 5.3.1 MAX-SMTI problems
First, we compare the solution quality of different algorithms in solving small-sized MAX-SMTI problems, including 60 configurations. The maximum number of iterations is set to 3,000 for TBLS and TBLS-E, and 50,000 for LTIU, AS, and MCS, to fully leverage the capabilities of the latter three local search algorithms.
@@
Figure 1(a) displays the number of singles produced by various algorithms in solving small-sized MAX-SMTI problems. Our TBLS algorithm achieves the highest number of wins and the lowest average number of singles across 60 configurations. TBLS-E closely follows, also demonstrating excellent performance with an average number of singles very similar to that of TBLS. Furthermore, TBLS-E stands out as the top performer measured by SECost, as illustrated in Figure 1(b). In almost all configurations, TBLS-E achieves the fairest matchings, significantly outperforming other algorithms.
Next, we assess the solution quality in addressing large-scale MAX-SMTI problems across 100 configurations, maintaining the same settings for the maximum number of iterations as used for small-sized MAX-SMTI problems. Due to its slow performance, LTIU is excluded from this evaluation. As the instances become harder, the performance gap between algorithms widens. Regarding the number of singles, TBLS remains the winner in most configurations, with approximately 0.2 fewer singles than the second-place TBLS-E and about 1.6 fewer singles than the third-place GSA2, as shown in Figure 2(a). As illustrated in Figure 2(b), TBLS-E consistently leads in most configurations, with an average SECost approximately 100 lower than MCS, which ranks second among the algorithms specifically designed for the MAX-SMTI problems.
@@
(a) Small-sized MAX-SMTI - execution times
(b) Large-sized MAX-SMTI - execution times
@@
It is noteworthy that the proposed TBLS-E algorithm outperforms both PDB and all other algorithms, achieving the lowest SECost in 95% of small-sized configurations and 78% of large-sized configurations. This improved performance can be attributed to the fact that, in addition to leveraging PDB as its base algorithm, TBLS-E introduces further constraints on tie-breaking adjustments to better balance satisfaction. These results support the effectiveness of TBLS-E in identifying more equitable matchings while maintaining matching sizes comparable to those achieved by its original algorithm TBLS.
Finally, we evaluate the execution times of different local search algorithms in solving small-sized and large-sized MAX-SMTI problems. For these comparisons, the maximum number of iterations is set to 3,000 for all the algorithms. Figure 3(a) illustrates the execution times for small-sized MAX-SMTI problems. Although all these algorithms, except LTIU, complete their searches within an average of 0.2 seconds, our TBLS algorithm performs the fastest in most configurations. This gap becomes more pronounced as the problem size increases, which is shown in Figure 3(b). For large-sized MAX-SMTI problems, TBLS and TBLS-E are the top performers in nearly all configurations. On average, the first-place TBLS and the second-place TBLS-E run about 2.5 and 1.2 times faster than the third-place MCS, respectively.
@@ 5.3.2 MAX-HRT problems
We begin by evaluating the solution quality of different algorithms for solving MAX-HRT problems of small and large sizes. The maximum number of iterations is still set to 3,000 for TBLS. To maximize the performance of other local search algorithms, the maximum number of iterations is set to 5,000 for both HR and MCA.
Figure 4(a) and Figure 4(b) illustrate the number of unassigned positions produced by various algorithms in addressing small-sized and large-sized MAX-HRT problems, respectively. Overall, TBLS outperforms all other algorithms in both problem scales, consistently yielding matchings with the fewest unassigned positions. Notably, for large-sized MAX-HRT problems, TBLS produces approximately 1.8 fewer unassigned positions than the second-place algorithm, HPA, and about 11.8 fewer unassigned positions than HR, the second-best among all local search algorithms.
@@
Additionally, the experimental results reveal that existing local search algorithms suffer a notable decline in performance with increasing problem size $n$. This phenomenon is more obvious for the MAX-HRT problem, as illustrated in Figure 4. One possible explanation is that as input size grows, the solution space expands substantially, and local optima become more difficult to escape. In contrast, our proposed algorithms, TBLS and TBLS-E, maintained robust performance on both problems with larger input sizes. This may be explained by their distinct neighborhood structures: TBLS and TBLS-E employ tie-breaking strategies, whereas other local search algorithms rely on removing blocking pairs. These findings suggest that tie-breaking based neighborhood structures facilitate more effective exploration of the solution space, thereby enhancing both solution quality and robustness, particularly in large-scale instances.
(a) Small-sized MAX-HRT - execution times
(b) Large-sized MAX-HRT - execution times
@@
After evaluating the solution quality, we shift our focus to analyzing execution time performance. The maximum number of iterations is set to 3,000 for all the algorithms. Figures 5(a) and Figure 5(b) show the execution times for small-sized and large-sized MAX-HRT problems, respectively. In both problem scales, TBLS emerges as the algorithm with the highest number of wins, indicating that it runs the fastest in a large number of configurations. In particular, for large-sized MAX-HRT problems, TBLS operates approximately 2.6 times faster than MCA and about 3.2 times faster than HR.
5.3.3 Scalability
In this section, we increase the input size $n$ from $1,000$ to $10,000$ to evaluate the scalability of our proposed algorithms in comparison with other local search methods. Since this experiment is intended to observe the trend in running time, the maximum number of iterations for each algorithm is set to $3,000$. We use the average number of singles or unassigned positions to assess the solution quality of the algorithm, as defined in Equation (1). If an algorithm fails to produce any stable matching for a given input size, no data point is shown for that algorithm at that size in the figure reporting the average number of singles or unassigned positions.
As shown in Figure 6(a), the AS and MCS algorithms are unable to produce stable matchings for the MAX-SMTI instances within the given number of iterations when the input size exceeds 2,000 and 3,000, respectively. Since our algorithms resolve all ties, they are guaranteed to find a stable matching regardless of the number of iterations or the instance size. Figure 6(b) depicts the relationship between the input size and the running time for solving MAX-SMTI problems. All four algorithms exhibit growth in time that is slightly faster than linear as the input size increases from 1,000 to 10,000. However, our proposed algorithms appear to have a more moderate growth rate in execution time. Compared with the other two local search algorithms, TBLS and TBLS-E achieve significantly better solution quality and substantially lower running times, demonstrating greater scalability.
(a) Number of singles
(b) Execution times
Figure 6: Scalability of four local search algorithms on MAX-SMTI problems
(a) Number of unassigned positions
(b) Execution times
Figure 7: Scalability of three local search algorithms on MAX-HRT problems
Figure 7 compares the scalability of three local search algorithms on instances of the MAX-HRT problem. It is worth noting that TBLS produces perfect matchings for nearly all instances across various input sizes, whereas HR yields an average of 37 unassigned positions, and MCA fails to find stable matchings for instances when $n$ exceeds 7,000. Moreover, as illustrated in Figure 7(b), TBLS consistently outperforms the other two algorithms in terms of running time. Based on the observed trends over the tested input sizes (from 1,000 to 10,000), there is no evidence suggesting that the growth rate of TBLS’s running time will surpass that of the other two algorithms as the input size increases.
@@ 6 Conclusion
This paper proposed a tie-breaking based local search algorithm, named TBLS, to solve the MAX-SMTI and MAX-HRT problems. We also introduced an equity-focused variant, TBLS-E, designed to find relatively fair matchings for the MAX-SMTI problem. Experimental results demonstrate that TBLS consistently outperforms other algorithms in achieving larger matching sizes for both the MAX-SMTI and MAX-HRT problems. Additionally, for the MAX-SMTI problem, TBLS-E yields matchings with lower sex equality costs while preserving matching sizes comparable to those produced by TBLS. Both TBLS and TBLS-E exhibit faster execution times than other local search algorithms when solving large-scale problems. The results of our scalability test suggest that both algorithms continue to perform efficiently with increasing problem size. In the future, we plan to incorporate neural networks and reinforcement learning techniques, such as G3DQN [24] and NeuRewriter [5], to improve the evaluation function and the refinement process.